skip to content
BeppoSAX U.S. Coordination Facility

BeppoSAX Status

List of events regarding the satellite/operations status extracted from the SAX-SDC News.

* 11 April 2002 End of Bepposax Operations -

Due to the poor and degrading spacecraft conditions and to the rapid orbital decay ASI has decided to terminate all in-orbit operations of the BeppoSAX mission at the end of April 2002. This date approximately agrees with the end of AO5 and (sadly) coincides with the 6th anniversary of the BeppoSAX launch. On April 30th the spacecraft will therefore be deactivated and put into its final configuration in preparation for re-entry accordingly to the existing international safety regulations. All BeppoSAX ground segment activities will then stop except for FOT production in order to ensure the completion and final validation of the data archive which will be hosted and supported at the ASI Science Data Center for the foreseeable future.

* 23 February 2002 Bepposax Status -

Since January 23 2002, BeppoSAX has progressively lost 4 of the 32 cells of Battery 1. This evolution was somehow expected since October 2001 when Battery 1 first displayed clear signs of irreversible ageing. The loss of these four cells (23,24,29,30) has severely degraded the battery performance with important consequences on the electrical and thermal stability of the spacecraft. Up to Febraury 13 a reduced instrument configuration was anyway supported (PDS, MECS, WFC1,WFC2) in an attempt to continue the AO5 programme as much as possible. However, after that date no scientific activities could be carried out since all the efforts at the Control Center had to be reserved to the rescue of BeppoSAX from a progressive thermal and electrical degrade. Strategies devised by the ground control to support Battery 1 ranged from the increase of the Battery K factor to the switch off of one Battery Discharge Regulator for as much as 15 minutes during the sun eclipse. The net outcome of these activities was the shift of a part of the power load on Battery 2 which is in a healthier condition than Battery 1. Since the loss of the fourth cell (February 23 2002) the spacecraft has not reached thermo-electrical stability yet, while also Battery 2 has started displaying fatigue signals at the end of the discharging phase. Work is still in progress at the OCC in order to re-establish the nominal temperature of 5 thermal loops.The case of a possible restart of some reduced scientific activity will be examined after the nominal thermo-elecrical values are attained. Battery 2 alone has not been designed to support the mission even in a minimal payload configuration. After the loss of Battery 1, Battery 2 will probably only allow the OCC to switch BeppoSAX definitevely off before collapse. At this time the general situation of the Electric Power System is to be considered critical with the possibility of further degradation that could easily lead to a sudden end of the mission.

* 22 October 2001 2nd EMS2 installation update -

On Friday 19 Oct 2001 at 17:10:52 BeppoSAX has successfully acquired the first scientific target using the new ESM2 'gyroless' pointing system: X1822-371. This pointing is controlled only by STRs troughout the orbit. After several hours of stable control the MECS, PDS and WFCs have been activated and started producing nominal scientific data. Preliminary analysis of the MECS images indicate a full compliance with the mission requirements.

* 1 October 2001 EMS2 installation update -

The ESM2 commissioning activities have officially started today. The interruption of AO5 observations is scheduled for Thursday 4th 2001 at noon time. The spacecraft will then be prepared for the loading of the new control mode and for the commissioning. Scientific activities are foreseen to restart on October 30 within the ESM2 scenario, pending the reliability and robustness of the new mode. As of today no AO5 TOO requests can be accepted. Only 'fast GRB repointings' will be considered up to next Thursday.

* 19 September 2001 EMS2 gyroless -

ESM2 (gyroless) attitude control mode installation schedule After the clear signs of degradation recently shown by the last surviving BeppoSAX GYRO, ASI has decided to speed up operations to install the ESM2 'gyroless' attitude control mode. The current schedule foresees a 3 week period of operative commissioning and ground segment reconfiguration. The planned AO5 scientific activity will be interrupted at the end of the first week of October to be resumed after October 29, pending the positive conclusion of the commissioning and the reliability of the ESM2 mode.

* 13 September 2001 Resumed operation -

BeppoSAX scientific operation resumed. Following a period of close monitoring of the GYRO6 behaviour BeppoSAX has resumed its scientific program. Preparations for the installation of the ESM2 "gyroless" poiting system are however ongoing. The plan of the ESM2 installation activities will be available shortly.

* 4 September 2001 Gyro status -

BeppoSAX Gyro6 update. On 21 and 28 August 2001 BeppoSAX suffered two losses of attitude which triggered a 'safe' mode (GSM). In both cases this was due to a sudden increase of the output from the last surviving gyro (GYRO6) with a dramatic change in the drift estimate. This is normally interpreted as a sign of the gyro degradation which leaves a slim residual lifetime. Since August 28 BeppoSAX has been put out of service in a safe configuration and the GYRO6 behaviour is monitored carefully. No recurrences of the gyro output increase has been since detected. Pending the stabilization of the gyro BeppoSAX will resume the planned scientific activity today however with an increased level of alert on the gyro measurements and operation in general.

* 3 May 1999 Resumed operation -

BeppoSAX has resumed normal science operations starting on May 1st at 06:30 UT

* 23 April 1999 Attitude status -

We are finishing the analysis of the accident which caused BeppoSAX to switchover its attitude control computer in 14th April. This analysis has taken us longer than we originally estimated. The cause of the switchover was a corruption in the onboard safe mode attitude control software. We are still analyzing the cause, and the methods to better monitor the onboard software status to avoid a similar problem in the future. We intend to return to the normal scientific observation program by Thursday 29th April.

* 16 April 1999 Attitude status -

On Wednsday 14 April at 07:30 GMT we had a switchover of the attitude control computer. We are still investigating the reason prior to returning to the scientific mission plan which we expect to do next week.

* 25 May 1998 WFC1 status -

Wide Field Camera 1 has returned to its normal observational activity two weeks ago (12/5/98) with its full operational and scientific performance intact. On Saturday 11th April WFC1 suddenly started to show a 3% instability on the high voltage of its position sensitive proportional counter. Immediately this was realized it was switched off. In collaboration with the personnel of SRU/SRON who constructed the WFC, the subsequent investigation showed that the high voltage was unstable even at values lower than nominal. A campaign of very small increments in voltage was started from well below the onset of the instability. This was done during the 10 minutes visibility periods in the orbits during the day when the consequences of each increment could be evaluated in real time. Occasional instability occurred but disappeared almost immediately. The last sign of instability occurred at 2960 volts from which we reached the nominal working voltage of the counter at 3727 volts with no further incident. Single orbit tests at reduced voltage were then performed without the problem returning. It was thus decided to return WFC1 to normal operation on 12/5/98. It is difficult to reach a firm conclusion about the nature of the problem. Now WFC1 has been functioning perfectly for about two weeks (or two hundred orbits) and its recovery looks permanent. We are all very content to be able to welcome WFC1 back to full participation in the BeppoSAX mission.

* 15 sep 1997 Operation Status -

On Friday 12 September scientific operations have been interrupted as a consequence of an on-board data corruption in a star sensor induced by a cosmic ray. Normal operations will restart later today.

* 9 sep 1997 Reccommencement of TOO Activity -

We are pleased to announce that from 9 September 1997 we are reccommencing our normal Target Of Opportunity activity. The new single gyro scientific mode software is performing very well.

* 14 Aug 1997 Science Operation Restarted -

The commissioning of the Single Gyro Scientific Mode software has now been succesfully completed. A first analysis of the data acquired reveals that the accuracy of the pointing is essentially the same as when three gyros where in use. Normal scientific activity has been restarted today.

* 11 Aug 1997 Sigle Gyro Pointing Software -

The commissioning of the Single Gyro Scientific Mode software has started on Aug 7th and is proceeding smootly. A few observations using the new pointing software have already been succesfully completed. We expect to resume routine observations within a few day

* 17 July 1997 BeppoSAX, satellite status -

The Single Gyro Scientific Mode software is now undergoing its final testing at Alenia Spazio, Turin. The commissioning of the new operational mode should commence on the satellite on Monday 28 July 1997. The commissioning is expected to take two weeks so that the normal scientific program of observations will restart by the 11th August.

* 6 June 1997 BeppoSAX, satellite status -

On thursday, 29/05/97, gyro 2 suddenly showed a reduction in sensitivity of a factor of two (in scale factor). This situation worsened over the next two days. At this point the gyro was switched-off for safety reasons.
The loss of gyro 2 leaves us with only two healthy gyro's (no.5 and no.6)- both of which are still performing well. We would need three healthy gyro's to continue the normal scientific mission until arrival of the single gyro and gyroless on-board attitude control software, and thus until then the satellite will operate in its gyroless safe mode.

We confirm that the single gyro software should be ready by the mid of July and the gyroless version at the beginning of October. As expected, the simulations of these modes demonstrate that the characteristics of the mission will remain very similar to those up to now.

For the period prior to the arrival of the single gyro mode we are working on a plan of scientific activity which foresees at least the following:
1 ) Gamma Ray Burst monitoring with the monitor (GRBM) in the phoswich detector system.
2 ) Operation of the wide field cameras (WFC) at least in the part of the orbit in sunlight where there is no danger that they could point towards the Sun or at Sco X-1 while observing.
With this plan we should at least be able to obtain the X-ray spectra of coincident bursts (seen simultaneously by the GRBM and WFC). If there are very strong known X-Ray sources in the WFC field of view the direction of the burst may be calculable. Also known x-ray source monitoring may be possible.

We will keep you informed of all developments.

* 23 May 1997 BeppoSAX Gyro Status -

In the evening of Tuesday 20 May 1997 Gyro-1 caused a fallback from a scientific pointing. Analysis of its output during the next 2.5 days has shown that its parameters (scale factor and drift rate) have changed and are evolving. It is evident that this gyro is no longer useful for scientific operations. As a consequence we have put the satellite into its gyro-less safe mode, and interrupted the scientific program. This was a suffered decision as if it was at all realistic to have continued using this gyro we would have made all efforts to do so.
We are now proceeding to the introduction of gyro-2 to replace gyro-1 in prospect to a return to the normal scientific program by Wednsday 28th May.
While we are preparing to return to normal operations we are still operating the Gamma-Ray burst monitor and the two Wide Field Cameras to look for coincident events. We will not however be able to give tempestive response on direction for follow-up observations as in the gyro-less safe mode the axes of the satellite follow the Earth's magnetic field, and the Sun, when available.

* 9 May 1997 BeppoSAX MECS Status -

In the afternoon of May 6 1997 the detector unit 1 of the MECS developed a fault. The investigation and on-board testing performed during the subsequent two days have localized the problem to the unit's gas cell high voltage supply. It is not currently possible to say if this unit can be recovered, but the investigation is still continuing. The other two detector units of the MECS (DU2 and DU3) are working normally. Normal scientific operations of the MECS using DU2 and DU3 is restarting today.

* 26-Feb-1997 BeppoSAX Gyro Status -

As a result of the several gyro problems which stopped the observation program from 31-Jan-1997 to 20-Feb-1997 a new series of pointing modes are in preparation. The current (nominal) pointing mode needs any three of the four gyros that are currently operational. The new modes need only one or no gyros in the control loop and are expected to retain the same scientific quality that can be achieved at present.

* 21 Feb 1997 Resumed Normal Operations -

Normal science operations have been resumed on Feb 20 at 10:00 GMT.

* 10 Feb 1997 BeppoSAX Gyro Status -

On 31 January 1997 one of the BeppoSAX gyros started to become unstable. This is the second gyro since the start of the mission that has given problems. There are 6 gyros on board and three are needed to continue the normal observation program. In consequence of this event it has been decided to temporarily interrupt the observation program and to commission a new safe mode for the satellite, called the Gyro-less Safe Mode. This safe mode was succesfully implemented on-board the week starting 3rd February. It has been decided to use this mode as baseline for the future and all procedures are currently being revised. Normal science operations are expected to be resumed within a few days.

* 20 Dec 1996 Resumed Normal Operations -

The problems encountered by the satellite on December 11 (see bullet dated Dec 17) have now been solved. Normal science operations have been resumed.

* 17 Dec 1996 Reset of the ACC-

At 17:30 GMT on Wednesday 11 December the satellite suffered a reset of the Attitude Control Computer which brought BeppoSAX into a completely safe full-back mode, as forseen, stopping nominal science operations.
Our investigations have shown that there is no problem in the hardware but the reason of the reset is still under investigation and we intend to recommence our normal observation activity as soon as the reason for the reset has been understood. At this moment BeppoSAX is continuing to monitor for gamma ray bursts with the PDS and WFC are being used to monitor for transient phenomena.

* 28 Oct 1996 On-board contingency -

Recently Beppo-SAX observations have been disturbed by an on-board contingency which has now been resolved and the normal scientific observations resumed.
On 16/10/1996 a persistent problem of syncronization between the attitude control computer and the data handling computer was detected. A software patch for the data handling computer was developed and tested to force the correct syncronization. This was sent up to the satellite on 23/10/1996. The patch was entirely successful, and after careful check-out the normal planned schedule of observation was restarted on the evening of 24/10/1996.
During contingency the observation time was not entirely lost as the NFI made background measurements and the WFC continued to monitor the Galactic Centre region and particularly GRS1915+105, the superluminal transient, which is currently active.

* 13 Aug 1996 SVP Observation -

SAX has successfully observed the following SVP observations: Cyg Field with WFC1, Cas A, SCO X-1 and 4U1626-67 with NFI, Cen X-3 with WFC2.
The Long Term Plan for Core and Guest Observers Program targets is in preparation. This plan takes into account the new constraint on the LECS which currently is operated only when the spacecraft is above dark earth. This constraint will be relaxed somewhat in the next few weeks.

* 24 Jun 1996 SAX first observation -

SAX has performed the first observation of a cosmic source.

On saturday 22 June Cygnus X-1, a bright blackhole candidate in our galaxy, was successfully observed with all NFI instruments and WFCs of SAX. Data analysis of this observation is in progress.
* 3-17 June 1996 Instruments commissing phase -

17 Jun 1996
The telescopes shutters have been successfully opened on saturday, June 15. The LECS's commissioning phase is going on and will finish tomorrow. The HPGSPC's commissioning phase is scheduled for June 19 and 20.

10 Jun 1996
The three MECS detectors have been swithched on at their nominal settings between June 6 and June 8th. The behaviour of all MECS instruments is nominal. Since the telescopes shutters are still closed the images obtained only include local background and calibration sources. A first M2 image is available. More detailed information is available from the
MECS team pages at IFCTR.

05 Jun 1996
A first WFC observation was carried out succesfully. The instrument is working nominally and the background intensity is as expected. The first detector image can be retrieved here

03 Jun 1996
The first SAX scientific instrument was successfully switched on. During days 29/5 - 1/6 the PDS switch-on procedure was successfully completed. This procedure was designed to switch on all the PDS subsystems, one at a time. The full switch on procedure took 22 orbits, during which the PDS was first operated at low values of HV, then at the nominal values. The PDS Particle Monitor is on since May 22. So now are already avalaible data on the particle fluxes during passages in the South Atlantic Anomaly for more than 10 days. Data are currently being analyzed. On saturday June 1st, PDS was switched on (all subsystems) for an entire SAX orbit (about 100 minutes). The measured background is consistent with the expected one. Data are being analyzed and in the next few days a first estimate of PDS background in coarse energy bands will be available. The high voltage verification activity on WFC will be performed 4-5th June.

* 1-28 May 1996 Commissioning Phase

28 May 1996
Gyro 2 has been excluded from the attitude control loop. Gyro 2 has shown an anomalous sporadic increase of approx a factor of 2 in its drift rate. This gyro is still considered by the program to be functional but was substituted (by gyro 6) for reasons of security. The reasons of the drift are under investigation.
A problem of increased background during star search mode in the star trackers, during sunlight, has been found. A modification is being implemented in the software of the star trackers to remove the problem. HPGSPC and PDS collimators have been successfully unlocked.
Phase D3, PMT (HV) check out, of PDS commissioning starts on May 29 and is expected to finish on June 1.

25 May 1996
The commissioning phase is proceeding regularly. The science instrumens HV will be switched on starting the beginning of June.

22 May 1996
The electronics of all payload instruments have been swithched on in stand-by mode. The PDS particle monitor is also active providing first information on the in-orbit background.
AOCS activities revealed an increase of the gyro 2 drift. Currently gyro 2 has been replaced with gyro 4 in the attitude control loop.

14 May 1996
All star trackers misalignments have been calibrated. The commissioning phase continues with the tuning of the AOCS system. Payload Testing is scheduled to start early next week.

08 May 1996
The Science Pointing Mode has been tested succesfully. Gyro's calibrations is proceeding as scheduled and will be completed tomorrow. At that point phase A and phase B of the commissioning phase will be completed.

06 May 1996
All on-board sub-systems are working properly. The commissioning phase is continuing as planned.

03 May 1996
The satellite is in the nominal configuration, on the nominal orbit and all sub-systems are properly performing. The Operation Control activities have been dedicated to the following main items: AOCS (Attitude & Orbit Control System) initialization in Default Pointing Mode (DPM), On Board Data Handling (OBDH) sub-system commissioning start, Telemetry Tracking & Control (TT&C) sub-system start.

02 May 1996
After the first 31 orbits all systems work properly. Later today OCC will start the calibration of magnetometers and gyroscopes.

* 30 Apr 1996

SAX has been successfully launched on Apr 30, 04:31 GMT, right at the beginning of the launch window.

If you have any questions concerning
BeppoSAX go to the Feedback form.

This file was last modified on Tuesday, 15-Jan-2008 11:42:19 EST